OpenStudy (anonymous):

A solution to the compound inequality 8 > 5 - 3x > -13 is provided below. Choose the statement that is true about the given solution. 8 > 5 - 3x and 5 - 3x > -13 Line 1 3 > -3x and -3x > -18 Line 2 -1< x and x < 6 Line 3 Please help me solve. (Without giving me the answer)

OpenStudy (amistre64):

work it with the same rules as last time, but you have 3 options to move about in

OpenStudy (amistre64):

most times its worked from middle out

OpenStudy (anonymous):

so 8 = 5 - 3x and 5 - 3x = -13?

OpenStudy (amistre64):

yes, that is a fine way to work it

OpenStudy (anonymous):

ok so first on 8 = 5 - 3x. -5 to both sides then divide 3?

OpenStudy (amistre64):

yes, but divide by -3 ... and we are dividing by a negative; flip the sign

OpenStudy (anonymous):

ok so -1 < x??

OpenStudy (amistre64):

very good

OpenStudy (anonymous):

ok and for 5 - 3x = -13. -5 to both sides divide -3.

OpenStudy (anonymous):

and get x < 6?

OpenStudy (amistre64):

yes, and say to yourself; divide by negative, flip the sign

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so in the equation there would be no mistake made.

OpenStudy (amistre64):

you did good -1 < x < 6

OpenStudy (anonymous):

ok thank u for the help. again :P

OpenStudy (amistre64):

youre welcome